# A Comprehensive Multi-Vocal Empirical Study of ML Cloud Service Misuses

### v1

@@ 5. Discussion and Recommendations
Towards new methodologies for ML service misuses detection. Our study establishes a foundation for developing automated detection and refactoring tools to enhance the quality of ML service-based systems. While some ML misuses can be easy to detect and refactor (e.g., “Non specification of early stopping criteria”, “Misinterpreting output”), others (e.g. “Ignoring fairness evaluation”, “Ignoring monitoring for data drift”) can be more complicated and require dedicated techniques to address. By cataloging prevalent ML service misuses and their occurrences in open-source projects, we provide a concrete basis for building tools that automatically identify and refactor these issues. Future research should focus on extending and refining this catalog, and developing solutions that help practitioners adhere to best practices, thereby improving the quality and maintainability of such systems.
@@ 6. Threats To Validity
We describe in this section the threats to the validity of our study.
@@ 6.1. Construct validity:
When considering the construct validity of our study, some threats must be acknowledged. One key concern is the method used to collect GitHub projects with identified misuses. Our scripts for project identification may have unintentionally excluded or included certain projects that were not entirely relevant to the scope of our study. To mitigate this risk, we conducted a manual review of the projects to ensure their relevance to our research on ML service misuses. This additional step helped refine the dataset and enhance its accuracy.
Furthermore, while our catalog of 20 misuses is based on a thorough review of both research and gray literature, there is always a possibility that some misuses were overlooked. Despite its comprehensiveness, the catalog may not capture every potential misuse in ML services. For instance, platforms such as Stack Overflow could offer valuable insights into other crucial misuses not yet addressed in our catalog. To address this, we are extending the catalog in a parallel independent study, incorporating contributions from such platforms to further strengthen its completeness and relevance in the field of ML service misuses.
@@ 6.2. Internal validity:
Our search query might not cover all terms related to ML service misuses and could miss important research resources. We accept these threats as our goal was not to provide an exhaustive list of all existing ML service misuses. As a mitigation solution, we (1) included in our search query the most important keywords related to the software quality of ML service-based systems, and (2) applied forward and backward snowballing to minimize the risk of missing important resources.
Furthermore, social desirability is a bias that leads our survey participants to agree with misuses. To minimize this threat, we did not offer any incentives for practitioners to participate in our survey. We also guaranteed the practitioners their anonymity and emphasized that all the reported information would be for research purposes only.
Another internal threat to the validity of our study is the practitioners’ interpretation of the identified misuses. To mitigate this, we provided a link to our online catalog within the survey, allowing participants to access detailed descriptions of each misuse. In addition, we categorized the misuses according to their respective stages within the ML development pipeline, ensuring a structured understanding. The catalog also includes real-world examples of misuse occurrences, along with proper references, to facilitate a more informed and accurate assessment of the misuses. We also extended invitations for interviews to practitioners to collect additional insights. However, we did not receive any responses.
One potential threat to internal validity is the manual analysis, which could introduce bias in the findings. To mitigate this, two evaluators (coauthors of this paper, with experience in machine learning and software engineering) independently analyzed the projects based on the same criteria. All discrepancies were resolved through discussion and consensus meetings. The agreement level between the evaluators was measured using Cohen’s Kappa and showed a strong level of consistency, reducing the risk of bias in the analysis.
@@ 6.3. External validity:
These threats addresses the generalizability of our findings. While we surveyed 50 practitioners, we highly believe that this sample size is reasonable considering the length of the survey and the challenge of identifying professionals with specific experience in ML cloud services. Importantly, the participants represent a diverse group of practitioners, all actively involved in ML service-based projects. This diversity, coupled with the hands-on experience of the participants, strengthens the reliability of our findings as it reflects the insights of subject matter experts.
@@
Our study focuses on a sample of a curated set of 377 GitHub projects, all written in Python. While this sample provides valuable insights into ML service misuses, it may not fully represent the broader landscape of ML service misuses. The limited scope in terms of both project numbers and programming language means that the findings might not be directly generalizable to other programming languages. Future research should on expanding this sample to include a more diverse range of projects across different programming languages to enhance the generalizability and applicability of the results.
@@
Another external threat pertains to the coverage of reasons and mitigation solutions for ML misuses. To mitigate this threat, we mentioned possible reasons and mitigation solutions in the survey based on our literature review and practical experience. Before disseminating the survey, we conducted a pilot study with three professionals in the field to check the wording and coverage of the proposed answers. Furthermore, we provided “Other” options in the survey, allowing participants to freely input additional reasons and mitigation solutions for ML service misuses. None of the 50 practitioners suggested any additional reasons or mitigation solutions, which indicates that the predefined options comprehensively captured the key factors related to these elements. Despite these efforts, we acknowledge that further research is needed to deepen our understanding of the underlying causes of ML service misuses in software systems and to develop more effective mitigation strategies.
@@ 7. Conclusion
This paper presented a comprehensive, multi-vocal empirical study of misuses of ML services. We propose a catalog of 20 ML service misuses based on a systematic literature review, a gray literature review, and an analysis of GitHub projects. We defined these misuses in detail, along with their impact on ML service-based systems and the stage at which they are more likely to occur. We validated our catalog through an online survey with ML practitioners, in which we measured the level of their agreement with our catalog and the frequency of encountering each misuse in practice. In addition, we reported, based on practitioners’ experience, the common reasons behind the presence of such misuses and the recommended solutions to mitigate them. Our study offers developers interested in ML cloud services a thorough understanding of the bad practices to avoid when developing ML service-based systems. Moreover, our study raises awareness of ML service misuses and encourages the involvement of various stakeholders, including cloud providers, open-source platforms, practitioners, and researchers, to mitigate them. In future work, we aim to develop a tool to automatically detect the presence of these misuses, enhancing the accuracy and speed of identification processes. We also intend to develop an automated refactoring solution to effectively remove such misuses, thereby improving code quality and maintainability.

### v2

@@ 6. Discussion and Recommendations
Towards new methodologies for detecting ML service misuses. Our study establishes a foundation for developing automated detection and refactoring tools to enhance the quality of ML service-based systems. While some misuses can be relatively easy to detect and refactor (e.g., “Non specification of early stopping criteria”, “Misinterpreting output”), others (e.g., “Ignoring fairness evaluation” or “Monitoring for data drift”) may require more sophisticated techniques. Each misuse in our catalog is tied to a specific ML development stage and is often reflected in concrete code patterns or configurations, making them amenable to static or dynamic analysis. For instance, misusing synchronous APIs or missing checkpointing logic can be identified through source code inspection, while deployment misconfigurations (e.g., lack of versioning or disabled auto-scaling) can be detected in infrastructure-as-code or configuration files. These characteristics open opportunities for building static analyzers, IDE plugins, or CI-integrated checkers to proactively detect and address ML service misuses. By cataloging these issues and linking them to observable artifacts, we provide a concrete basis for tooling that supports proactive quality assurance. Moreover, future research should also invest in extending and refining the proposed catalog, formalizing misuse patterns, and developing practical tools to help practitioners adhere to best practices and improve maintainability.
Towards more awareness of ML service misuses. The positive correlation between practitioners’ experience levels and the frequency of encountering ML service misuses highlights the importance of knowledge transfer from experienced to less experienced practitioners. This emphasizes the need to raise awareness about best practices in ML service usage and improve the overall quality of ML service-based systems. This carries significant implications for several stakeholders in the ML ecosystem, including software and ML engineers, ML cloud providers, open-source platforms like GitHub, and researchers.
  • For practitioners: Our catalog of ML service misuses offers valuable insights into common pitfalls and misconfigurations that can compromise the reliability of ML service-powered systems. With this comprehensive reference, practitioners can better understand potential risks and proactively design and integrate ML services that align with best practices. For example, software engineers, who often serve as the bridge between model development and production environments, can leverage this catalog to mitigate risks related to fairness, scalability, and performance.
  • For ML cloud service providers: Our study highlights the importance of developing tools and platforms to facilitate the detection and prevention of ML service misuse. Some misuses, such as “Overwriting existing ML APIs without versioning”), are relatively straightforward and could be automatically detected. Building such tools would help ensure users benefit from more transparent, scalable, and high-performance ML services.
  • For open-source platforms: Platforms such as GitHub and other developer communities can play a critical role in promoting collaboration to identify and address these misuses. They can support the creation of shared resources that enhance the performance of ML service-based systems. For instance, GitHub could integrate tools that detect common ML service misuses during pull requests or commits, flagging potential issues early in the development lifecycle.
  • For researchers: Our catalog highlights several underexplored areas in ML service misuse that warrant deeper investigation, such as the underlying rationale behind practitioner decisions and the impact of emerging ML cloud technologies on misuse patterns. Researchers can build upon our findings to develop educational resources and design frameworks that guide practitioners toward best practices. Future research should also focus on the development of automated detection and refactoring tools specifically targeting ML service misuses.
Overall, our findings trigger an alarm towards investing in more robust tools, policies, and guidelines to safeguard the integrity, reliability, and societal impact of ML cloud services in particular, and ML-based systems in general.
@@ 7. Threats To Validity
This section outlines the threats to the validity of our findings.
@@ 7.1. Construct validity:
Construct threats to the validity concern the relationship between the theory and the observations made. Regarding our findings, one key concern is the method used to collect GitHub projects with identified misuses. Our scripts for project identification may have unintentionally excluded or included certain projects that were not entirely relevant to the scope of our study. To mitigate this risk, we conducted a manual review of the projects to ensure their relevance to our research on ML service misuses. This additional step helped refine the dataset and enhance its accuracy.
Furthermore, while our catalog of 20 misuses is based on a thorough review of both research and gray literature, there is always a possibility that some misuses were overlooked. Despite being thorough, the catalog may not capture every potential misuse in ML services. For instance, platforms such as Stack Overflow could offer additional valuable insights into other crucial misuses not yet covered in our catalog. To address this, we aim to extend the catalog in a parallel independent study, incorporating contributions from such platforms to further strengthen its completeness and relevance in the context of ML service misuses.
@@ 7.2. Internal validity:
Internal threats to validity are concerned with the causal relationship between treatment and outcome. Our search query might not cover all terms related to ML service misuses and could miss important research and gray literature. We accept these threats as our goal was not to provide an exhaustive list of all existing ML service misuses. As a mitigation solution, we (1) included in our search query the most important keywords related to the software quality of ML service-based systems, and (2) applied forward and backward snowballing to minimize the risk of missing important resources.
Furthermore, social desirability is a bias that leads our survey participants to agree with misuses. To minimize this threat, we did not offer any incentives for practitioners to participate in our survey. We also guaranteed the practitioners their anonymity and emphasized that all the reported information would be for research purposes only. Another internal threat to the validity of our findings is the practitioners’ interpretation of the identified misuses. To mitigate this, we provided a link to our online catalog within the survey, allowing participants to access detailed descriptions of each misuse. In addition, we categorized the misuses according to their respective stages within the ML development pipeline, ensuring a structured understanding. The catalog also includes real-world examples of misuse occurrences, along with proper references, to facilitate a more informed and accurate assessment of the misuses.
To complement the quantitative findings, our survey included open-ended questions aimed at capturing qualitative insights from practitioners regarding the misuse and best practices of ML services. While several participants responded to the open-ended questions, their answers were brief and lacked depth (e.g., expressing appreciation or thanking the researchers), making it difficult to conduct a meaningful thematic analysis. To mitigate this threat, we excluded these responses from our analysis and instead relied on qualitative evidence from our gray literature review (e.g., online resources, documentation, and prior research papers). We further extended invitations for interviews to practitioners to collect additional insights. However, we did not receive any responses. This aligns with known challenges in software engineering research, where recruiting practitioners for interviews has been widely reported as difficult  (ghaleb2022studying, ; kokinda2023under, ). We acknowledge that interview-based data could have provided greater depth and consider this an avenue for future work.
One potential threat to internal validity is manual analysis, which could introduce bias in the findings. Moreover, the notion of what is an ML service “misuse” may be open to interpretation due to the lack of a universally formalized definition in the literature. To mitigate this, two evaluators (co-authors of this paper, with experience in machine learning and software engineering) independently analyzed the projects based on the same criteria. All discrepancies were resolved through discussion and consensus meetings. The agreement level between the evaluators was measured using Cohen’s kappa and showed a strong level of consistency, reducing the risk of human subjectivity and error in the analysis.
@@ 7.3. External validity:
These threats address the generalizability of our findings. Our study focused on a sample of a curated set of 377 GitHub projects, all written in Python. While this sample provides valuable insights into ML service misuses, it may not fully represent the broader landscape of ML service misuses. The limited scope in terms of both project numbers and programming language means that the findings might not be directly generalizable to other programming languages. Future research should focus on expanding this sample to include a more diverse range of projects across different programming languages to enhance the generalizability and applicability of the results.
Moreover, some misuses in our catalog, such as “Not using batch API for data processing” and “Ignoring fairness evaluation,” may not apply in all contexts. For example, batch API misuse is relevant only when batch processing is required, and fairness evaluation mainly concerns datasets with protected attributes. However, fairness assessment remains important, as biases can also emerge from complex feature interactions or evolving populations. To address this, we clarified in the catalog the conditions and contexts in which these misuses apply. For example, we distinguished between batch and real-time streaming data when identifying batch API misuse to reduce the risk of misclassifications and misunderstandings. While our catalog is specifically designed around cloud-based ML services, several of the misuses identified are also relevant to ML systems more broadly. However, the impact of the identified misuses in cloud environments is more pronounced and leads to notable effects, such as increased latency, higher costs, and degraded performance. These impacts may be different or less severe in non-cloud or traditional ML environments (yao2017complexity, ).
Though our survey included 50 practitioners (a 12% response rate), we believe this sample is somewhat reasonable, as achieving high response rates is challenging in this kind of studies (ghazi2018survey, ), especially in our context given the survey length and the difficulty of identifying professionals with specific experience in ML cloud services. Importantly, the participants represent a diverse group of practitioners, all actively involved in ML service-based projects. This diversity, coupled with the hands-on experience of the participants, strengthens the reliability of our findings as it reflects the insights of subject matter experts. Nevertheless, we acknowledge that the final response rate (12%) and the small number of respondents may introduce self-selection bias and limit the statistical generalizability of our results to the broader ML practitioner population. To mitigate this risk, we reached out to participants through diverse channels (e.g., professional networks and email), ensured representation of different roles and experience levels, and limited outreach to no more than five individuals per company to avoid over-representation of any single organizational perspective.
@@
Another external threat pertains to the coverage of reasons and mitigation solutions for ML misuses. To mitigate this threat, we mentioned possible reasons and mitigation solutions in the survey based on our review of the literature and practical experience. Before disseminating the survey, we conducted a pilot study with three professionals in the field to verify the language and coverage of the proposed answers. Furthermore, we provided an “Other” option in the survey, allowing participants to freely input additional open-ended reasons and mitigation solutions for ML service misuses. Yet, none of the 50 practitioners suggested additional reasons or mitigation solutions, indicating that the predefined options comprehensively captured the key factors related to these elements. Despite these efforts, we acknowledge that further research is needed to deepen our understanding of the underlying causes of ML service misuses in software systems and to develop more effective mitigation strategies.
@@ 8. Conclusion
This paper presented a comprehensive, multi-vocal empirical study of misuses of ML cloud services. We propose a catalog of 20 ML service misuses based on a research literature review, a gray literature review, and an analysis of GitHub projects. We defined these misuses in detail, along with their impact on ML service-based systems and the stage at which they are more likely to occur. We validated our catalog through an online survey with ML practitioners, in which we measured the level of their agreement with our catalog and the frequency of encountering each misuse in practice. In addition, we reported, based on practitioners’ experience, the common reasons behind the presence of such misuses and the recommended solutions to mitigate them. Our study offers developers interested in ML cloud services a thorough understanding of the bad practices to avoid when developing ML service-based systems. Moreover, our study raises awareness of ML service misuses and encourages the involvement of various stakeholders, including cloud providers, open-source platforms, practitioners, and researchers, to mitigate them. In future work, we aim to explore additional ML service misuses from sources such as Stack Overflow to expand our catalog, as well as develop a tool to automatically detect the presence of these misuses, enhancing the accuracy and speed of identification processes. We also intend to create an automated refactoring solution to effectively remove such misuses, thereby improving code quality and maintainability.